# SoTA with Less: MCTS-Guided Sample Selection for Data-Efficient Visual Reasoning Self-Improvement

### v2

@@ Abstract
We introduce ThinkLite-VL, a family of visual reasoning models that achieve state-of-the-art (SoTA) performance using an order of magnitude fewer training samples, relying purely on reinforcement fine-tuning (RFT) self-improvement without any knowledge distillation. Our central insight is that sample difficulty critically influences RFT effectiveness: appropriately challenging examples can drive substantial reasoning improvements, even in low-data regimes. However, quantifying sample difficulty in a reliable and scalable manner remains non-trivial. To address this, we repurpose Monte Carlo Tree Search (MCTS) to measure sample difficulty via the number of reasoning iterations a vision-language model (VLM) requires to solve each instance. This MCTS-based selection procedure identifies samples that induce deeper reasoning while remaining solvable, allowing us to filter a high-quality subset from 70k open-source examples spanning math, natural image understanding, and chart comprehension. Using this approach, we select just 11k challenging samples for RFT on Qwen2.5-VL-7B-Instruct and 7.5k samples for Qwen2.5-VL-72B-Instruct. The resulting models, ThinkLite-VL-7B and ThinkLite-VL-72B, significantly outperform their respective base models across eight visual reasoning benchmarks. In particular, ThinkLite-VL-7B improves the average performance of Qwen2.5-VL-7B-Instruct by 7% and surpasses all existing 7B-level models, as well as much larger models such as GPT-4o, O1 and Qwen2.5-VL-72B, achieving a new SoTA score of 75.1 on MathVista. ThinkLite-VL-72B further advances the SoTA frontier, achieving an accuracy of 79.7 on MathVista and an average benchmark improvement of 4.42 over the open-source SOTA. These results demonstrate that MCTS-guided difficulty filtering provides a scalable and effective path toward data-efficient self-improvement in multimodal reasoning. Our code, data, and model are available at https://github.com/si0wang/ThinkLite-VL.
@@ 1 Introduction
Large language models (LLMs) have demonstrated strong capabilities in solving complex reasoning tasks—such as mathematics and coding—by leveraging chain-of-thought prompting and reflection mechanisms [24, 36]. Recent work [16] highlights the critical role of reinforcement fine-tuning (RFT) in further enhancing reasoning performance. Remarkably, these improvements can be achieved purely via RFT, even without post-training supervised fine-tuning (SFT).
@@
However, despite the success of RFT in LLMs, its impact on vision-language models (VLMs) has been less pronounced. A likely cause is the inherent modality gap: VLMs are pretrained on text-heavy objectives, while post-training tasks demand multimodal reasoning. Recent efforts [22, 12, 55, 84] have addressed this by incorporating knowledge distillation and supervised format alignment before RFT. While effective, these pipelines are cumbersome, and fundamentally limit the capacity for models to improve via self-training alone.
@@
    Strong empirical gains. We demonstrate that ThinkLite-VL-7B and ThinkLite-VL-72B outperform strong baselines and existing SoTA models across eight VLM benchmarks. Notably, ThinkLite-VL-7B improves the average performance of its base model by 7%, and achieves a new SoTA score of 75.1% on MathVista—surpassing larger models such as GPT-4o, O1 and Qwen2.5-VL-72B. ThinkLite-VL-72B further advances this with a MathVista score of 79.7%.
@@
    Open-source release. We will release the full ThinkLite-VL model family, including both ThinkLite-VL-7B and ThinkLite-VL-72B, and MCTS-filtered training sets for both Qwen2.5-VL-7B and Qwen2.5-VL-72B to support future research in multimodal reasoning.
@@ 2 Related work
Simulating human-like thinking processes through intermediate reasoning steps has significantly improved the performance of large language models (LLMs) on tasks that require reasoning [24]. One family of methods focuses on explicitly controlling the structure or format of the model’s outputs, such as by applying Chain-of-Thought (CoT) prompting [78] and Self-Consistency [77]. Related lines of work include more elaborate reasoning strategies like Tree of Thoughts [87] or Graph of Thoughts [4]. Additionally, some approaches involve supervised fine-tuning (SFT) on curated datasets with reasoning annotations [52, 88]. Researchers have also explored process reward models (PRMs) that encourage systematic thought processes [35, 66, 70, 28, 98, 47]. Others incorporate search techniques, including Monte Carlo Tree Search (MCTS) or beam search, to refine or verify reasoning paths [80, 81, 5, 15, 19, 73]. Recently, large-scale RL with outcome-based reward functions has been leveraged [16] to elicit powerful reasoning capabilities in LLMs. Unlike prior uses of MCTS at inference time [80, 81, 15], we employ MCTS during training to assess sample difficulty and curate a high-impact training subset for RFT. We focus on how to use large-scale RL to enhance the reasoning ability of VLMs.
@@
Vision language models [1, 68, 38, 23, 37, 3, 10, 64, 31, 85] can perform vision tasks using language given visual input through vision encoders like [57, 94, 65]. These models demonstrate comprehensive multimodal capabilities across various scenarios [93, 39, 91, 49, 18, 92, 20, 32] and exhibit reasoning capabilities to some extent [43, 76, 41, 96, 69]. Inspired by the success of reasoning in LLMs, researchers have sought to improve the reasoning capabilities of VLMs. For instance, CoT prompting is applied to VLMs [97, 51, 46, 11, 100, 21] and some papers create multimodal datasets [86, 83, 61, 99, 12, 22, 17, 63], using SFT for knowledge distillation to improve reasoning abilities. Some prior works have also explored improving VLM performance through self-improvement strategies [102, 72, 75, 13]. More recently, RL training has emerged as a promising approach to further strengthen the reasoning capabilities of VLMs [12, 22, 50, 82]. While recent works explore SFT and RL  [12, 22] for VLM reasoning, efficiently utilizing training data and avoiding costly knowledge distillation remains a challenge. In contrast, ThinkLite-VL eliminates the need for SFT or distillation entirely and achieves SoTA performance using just 11k (7B) and 7.5k (72B) samples—an order of magnitude less than prior work. Specifically, we propose a novel approach using MCTS to filter for high-quality training instances based on the difficulty level. We then directly apply RL training to enhance reasoning on this curated data, demonstrating strong performance without requiring any SFT stage.
@@
Data filtration aims to identify and retain high-quality, diverse, and task-relevant data while discarding noisy or redundant information to optimize training efficiency and generalization performance. It is important for the pretraining phase [14, 29, 79, 58, 54, 2, 95, 67, 56] and instruction tuning phase [34, 33, 7, 9, 38, 103, 89] of both LLMs and VLMs. In this paper, we specifically focus on filtering training instances to curate data optimally for efficient downstream RL training to improve the reasoning capabilities of VLMs. A concurrent work, MM-Eureka [50], also investigates the impact of data filtration on RFT. While MM-Eureka [50] filters samples based on zero-shot accuracy, our MCTS-based method provides a more expressive and fine-grained estimate of sample difficulty, capturing both solved and unsolved-but-informative cases. Importantly, our findings reveal that samples requiring extended reasoning—even when not solved by the model—can be highly beneficial during RFT.
@@ 3.1 Data Collection
We collect a total of 70k datas from widely used open-source training datasets as our initial training set, covering three category: multimodel mathematical reasoning (Geometry3K [42], GeoQA [6], Geos [60]), natural image understanding (FigureQA [25], ScienceQA [43], OK-VQA [48]), and chart understanding (IconQA [45], TabMWP [44]). For FigureQA and IconQA, due to the large size of their original training sets, we only randomly sample 10k data points from each as our training set. The overall data distribution is shown in Figure 3. Each training sample is organized into the following format: (Image, id, Prompt, Answer).
@@ 3.2 MCTS-based Sample Selection
In our work, the collected data primarily originates from commonly used pretraining datasets for existing VLMs, which makes the model susceptible to overfitting on certain samples. Inspired by recent successes of data filtration in LLM SFT [53, 88] and conventional reinforcement learning [59, 74], we propose a MCTS-based sample selection mechanism. This approach leverages the VLM’s own iterative reasoning process, using the number of iterations required to reach the correct answer as a metric to assess the difficulty of each data sample. Consequently, we can selectively filter for those samples that are more challenging for the model during RL training, rather than using the entire dataset.
@@ 3.3 Visual Reasoning Training
LLaVA-Cot-11B [83] GPT-4o 100k
Mulberry-7B [86] GPT-4o, Qwen2-VL-72B 260k
Vision-R1-7B [22] Deepseek-R1 200k + 10k
OpenVLThinker-7B [12] DeepSeek-R1-Distill-Qwen-14B 59.2k
MM-EUREKA-Qwen-7B [50] 15k
@@
After conducting MCTS-based sample selection and obtaining a filtered set of high-quality training data (11k for 7B and 7.5k for 72B), we then perform RL fine-tuning on the Qwen2.5-VL models using these selected data. Specifically, we employ Group Relative Policy Optimization (GRPO) loss function proposed by [62] for training, with the objective defined as follows:
@@ 4.1 Benchmark Evaluation
• For 7b-level VLMs, we use Qwen2.5-VL-7B-Instruct as the base model and perform RFT on the 11k high-quality data obtained through MCTS-based filtration, resulting in our reasoning model, named ThinkLite-VL-7B. We conduct training using Easy-R1 [101] code base and set GRPO rollout number as 32. Our main baselines are as follows: (1) Qwen2.5-VL-7B-Instruct [3], serving as our base model; (2) ThinkLite-VL-Random11k, trained using RFT on a randomly sampled subset of 11k instances from the full 70k dataset. Besides, we report the performance of several recent general and reasoning VLMs for comparison, including general opensourced models LLaVA-Onevision-7B [30] and InternVL2.5-8B [10], the SFT-based reasoning models LLaVA-Cot-11B [83] and Mulberry-7B [86], as well as the RFT-based reasoning models Vision-R1 [22], MM-Eureka-Qwen-7B [50], and OpenVLThinker-7B [12].
• For 72B-level VLMs, we use Qwen2.5-VL-72B-Instruct as the base model. We perform RFT on the 7.5k high-quality data obtained by Qwen2.5-VL-72B-Instruct through MCTS-based filtration and get 72B reasoning model ThinkLite-VL-72B. The 72B-level baselines include: (1) our base model Qwen2.5-VL-72B-Instruct [3]; (2) two opensourced general VLMs LLaVA-Onevision-72B [30] and InternVL2.5-78B [10]; (3) one opensouced reasoning model QvQ-72B [71]; (4) ThinkLite-VL-Random7.5k, trained using RFT on 7.5k randomly selected samples from the full 70k dataset. We also include proprietary models as performance references which include OpenAI-GPT-4o and OpenAI-o1. For all models, we use 8$\times$80G A100 GPUs for model training and evaluation.
@@
We select eight widely used VLM benchmarks for evaluation, namely MathVista [41], MathVison [69], MathVerse [96], MMMU [93], MMStar [8], MMBench [40], MMVet [91], and AI2D [26]. Among them, MathVista, MathVison, and MathVerse are widely used in VLM research to evaluate mathematical reasoning capabilities, while MMVet also includes a significant number of mathematical reasoning tasks. In contrast, MMMU, MMStar, MMBench, and AI2D are primarily utilized to assess VLM’s visual perception reasoning and scientific reasoning abilities.
Table 2: Comparison of different VLMs on 8 widely used visual benchmarks. Our model achieves SoTA performance at both 7B level and 72B level on 6 benchmarks and reaches a SoTA performance of 79.7 on MathVista among all VLMs. On average, our model improves performance by 7.5% and 6.5% compared with our base models Qwen2.5-VL-7B-Instruct and Qwen2.5-VL-72B-Instruct. We do not evaluate Mulberry-7B on MathVision because Mulberry-7B uses MathVision as training dataset. We evaluate all models with same code using vLLM [27] inference. For reasoning models, we use thinking templates provided in their codebase to generate thoughts and get the final answer.

### v3

@@ Abstract
We introduce ThinkLite-VL, a family of visual reasoning models that achieve state-of-the-art (SoTA) performance using an order of magnitude fewer training samples, relying purely on reinforcement fine-tuning (RFT) self-improvement without any knowledge distillation. Our central insight is that sample difficulty critically influences RFT effectiveness: appropriately challenging examples can drive substantial reasoning improvements, even in low-data regimes. However, quantifying sample difficulty in a reliable and scalable manner remains non-trivial. To address this, we repurpose Monte Carlo Tree Search (MCTS) to measure sample difficulty via the number of reasoning iterations a vision-language model (VLM) requires to solve each instance. This MCTS-based selection procedure identifies samples that induce deeper reasoning while remaining solvable, allowing us to filter a high-quality subset from 70k open-source examples spanning math, natural image understanding, and chart comprehension. Using this approach, we select just 11k challenging samples for RFT on Qwen2.5-VL-7B-Instruct and 7.5k samples for Qwen2.5-VL-72B-Instruct. The resulting models, ThinkLite-VL-7B and ThinkLite-VL-72B, significantly outperform their respective base models across eight visual reasoning benchmarks. In particular, ThinkLite-VL-7B improves the average performance of Qwen2.5-VL-7B-Instruct by 7% and surpasses all existing 7B-level models, as well as much larger models such as GPT-4o, O1 and Qwen2.5-VL-72B, achieving a new SoTA score of 75.1 on MathVista. ThinkLite-VL-72B further advances the SoTA frontier, achieving an accuracy of 79.7 on MathVista and an average benchmark improvement of 4.42 over the open-source SOTA. These results demonstrate that MCTS-guided difficulty filtering provides a scalable and effective path toward data-efficient self-improvement in multimodal reasoning.
= Date: May 30, 2025
[Uncaptioned image] Code Repository: https://github.com/si0wang/ThinkLite-VL
= Model Weights: https://huggingface.co/collections/russwang/thinklite-vl
= Datasets: https://huggingface.co/collections/russwang/thinklite-vl
= Contact: xywang@umd.edu
@@ 1 Introduction
Large language models (LLMs) have demonstrated strong capabilities in solving complex reasoning tasks—such as mathematics and coding—by leveraging chain-of-thought prompting and reflection mechanisms (Jaech et al., 2024, Liu et al., 2024a). Recent work (Guo et al., 2025) highlights the critical role of reinforcement fine-tuning (RFT) in further enhancing reasoning performance. Remarkably, these improvements can be achieved purely via RFT, even without post-training supervised fine-tuning (SFT).
@@
However, despite the success of RFT in LLMs, its impact on vision-language models (VLMs) has been less pronounced. A likely cause is the inherent modality gap: VLMs are pretrained on text-heavy objectives, while post-training tasks demand multimodal reasoning. Recent efforts (Huang et al., 2025, Deng et al., 2025, Peng et al., 2025, Yang et al., 2025) have addressed this by incorporating knowledge distillation and supervised format alignment before RFT. While effective, these pipelines are cumbersome, and fundamentally limit the capacity for models to improve via self-training alone.
@@
    Strong empirical gains. We demonstrate that ThinkLite-VL-7B and ThinkLite-VL-72B outperform strong baselines and existing SoTA models across eight VLM benchmarks. Notably, ThinkLite-VL-7B improves the average performance of its base model by 7%, and achieves a new SoTA score of 75.1 on MathVista—surpassing larger models such as GPT-4o, O1 and Qwen2.5-VL-72B. ThinkLite-VL-72B further advances this with a MathVista score of 79.7.
@@
    Open-source release. We release the full ThinkLite-VL model family, including both ThinkLite-VL-7B and ThinkLite-VL-72B, and MCTS-filtered training sets for both Qwen2.5-VL-7B and Qwen2.5-VL-72B to support future research in multimodal reasoning.
@@ 2 Related work
Simulating human-like thinking processes through intermediate reasoning steps has significantly improved the performance of large language models (LLMs) on tasks that require reasoning (Jaech et al., 2024). One family of methods focuses on explicitly controlling the structure or format of the model’s outputs, such as by applying Chain-of-Thought (CoT) prompting (Wei et al., 2022) and Self-Consistency (Wang et al., 2022b). Related lines of work include more elaborate reasoning strategies like Tree of Thoughts (Yao et al., 2023) or Graph of Thoughts (Besta et al., 2024). Additionally, some approaches involve supervised fine-tuning (SFT) on curated datasets with reasoning annotations (Muennighoff et al., 2025a, Ye et al., 2025). Researchers have also explored process reward models (PRMs) that encourage systematic thought processes (Lightman et al., 2023, Uesato et al., 2022, Wang et al., 2023b, Lai et al., 2024, Zhang et al., 2025, Luo et al., 2024). Others incorporate search techniques, including Monte Carlo Tree Search (MCTS) or beam search, to refine or verify reasoning paths (Xie et al., 2024, Xin et al., 2024, Chen et al., 2024a, Gao et al., 2024, Hao et al., 2023, Wang et al., 2024d). Recently, large-scale RL with outcome-based reward functions has been leveraged (Guo et al., 2025) to elicit powerful reasoning capabilities in LLMs. Unlike prior uses of MCTS at inference time (Xie et al., 2024, Xin et al., 2024, Gao et al., 2024), we employ MCTS during training to assess sample difficulty and curate a high-impact training subset for RFT. We focus on how to use large-scale RL to enhance the reasoning ability of VLMs.
@@
Vision language models (202, 2023, Wang et al., 2022a, Liu et al., 2023, Hurst et al., 2024, Liu et al., 2024b, Bai et al., 2025, Chen et al., 2024e, Tong et al., 2024, Li et al., 2024b, Yang et al., 2023) can perform vision tasks using language given visual input through vision encoders like (Radford et al., 2021, Zhai et al., 2023, Tschannen et al., 2025). These models demonstrate comprehensive multimodal capabilities across various scenarios (Yue et al., 2024, Liu et al., 2024d, Yu et al., 2024b, Masry et al., 2022, Gurari et al., 2018, Yu et al., 2024c, Hao et al., 2025, Li et al., 2025) and exhibit reasoning capabilities to some extent (Lu et al., 2022a, Wang et al., 2024f, Lu et al., 2024, Zhang et al., 2024a, Wang et al., 2024a). Inspired by the success of reasoning in LLMs, researchers have sought to improve the reasoning capabilities of VLMs. For instance, CoT prompting is applied to VLMs (Zhang et al., 2024b, Mitra et al., 2024, Luan et al., 2024, Chen et al., 2023, Zheng et al., 2023, Hu et al., 2024) and some papers create multimodal datasets (Yao et al., 2024, Xu et al., 2025, Shao et al., 2024a, Zhang et al., 2023b, Deng et al., 2025, Huang et al., 2025, Guo et al., 2024, Thawakar et al., 2025), using SFT for knowledge distillation to improve reasoning abilities. Some prior works have also explored improving VLM performance through self-improvement strategies (Zhou et al., 2024, Wang et al., 2024c, e, Deng et al., 2024). More recently, RL training has emerged as a promising approach to further strengthen the reasoning capabilities of VLMs (Deng et al., 2025, Huang et al., 2025, Meng et al., 2025, Xiong et al., 2024). While recent works explore SFT and RL  (Deng et al., 2025, Huang et al., 2025) for VLM reasoning, efficiently utilizing training data and avoiding costly knowledge distillation remains a challenge. In contrast, ThinkLite-VL eliminates the need for SFT or distillation entirely and achieves SoTA performance using just 11k (7B) and 7.5k (72B) samples—an order of magnitude less than prior work. Specifically, we propose a novel approach using MCTS to filter for high-quality training instances based on the difficulty level. We then directly apply RL training to enhance reasoning on this curated data, demonstrating strong performance without requiring any SFT stage.
@@
Data filtration aims to identify and retain high-quality, diverse, and task-relevant data while discarding noisy or redundant information to optimize training efficiency and generalization performance. It is important for the pretraining phase (Gao et al., 2020, Lee et al., 2021, Xie et al., 2023, Ruis et al., 2024, Penedo et al., 2024, Alayrac et al., 2022, Zhang et al., 2023a, Wang et al., 2023a, Radenovic et al., 2023) and instruction tuning phase (Li et al., 2023, 2024c, Chen et al., 2024b, d, Liu et al., 2023, Zhu et al., 2023, Yu et al., 2024a) of both LLMs and VLMs. In this paper, we specifically focus on filtering training instances to curate data optimally for efficient downstream RL training to improve the reasoning capabilities of VLMs. A concurrent work, MM-Eureka (Meng et al., 2025), also investigates the impact of data filtration on RFT. While MM-Eureka (Meng et al., 2025) filters samples based on zero-shot accuracy, our MCTS-based method provides a more expressive and fine-grained estimate of sample difficulty, capturing both solved and unsolved-but-informative cases. Importantly, our findings reveal that samples requiring extended reasoning—even when not solved by the model—can be highly beneficial during RFT.
@@ 3.1 Data Collection
We collect a total of 70k datas from widely used open-source training datasets as our initial training set, covering three category: multimodel mathematical reasoning (Geometry3K (Lu et al., 2021), GeoQA (Chen et al., 2022), Geos (Seo et al., 2015)), natural image understanding (FigureQA (Kahou et al., 2018), ScienceQA (Lu et al., 2022a), OK-VQA (Marino et al., 2019)), and chart understanding (IconQA (Lu et al., 2022b), TabMWP (Lu et al., 2023)). For FigureQA and IconQA, due to the large size of their original training sets, we only randomly sample 10k data points from each as our training set. The overall data distribution is shown in Figure 3. Each training sample is organized into the following format: (Image, id, Prompt, Answer).
@@ 3.2 MCTS-based Sample Selection
In our work, the collected data primarily originates from commonly used pretraining datasets for existing VLMs, which makes the model susceptible to overfitting on certain samples. Inspired by recent successes of data filtration in LLM SFT (Muennighoff et al., 2025b, Ye et al., 2025) and conventional reinforcement learning (Schaul et al., 2016, Wang et al., 2023c), we propose a MCTS-based sample selection mechanism. This approach leverages the VLM’s own iterative reasoning process, using the number of iterations required to reach the correct answer as a metric to assess the difficulty of each data sample. Consequently, we can selectively filter for those samples that are more challenging for the model during RL training, rather than using the entire dataset.
@@ 3.3 Visual Reasoning Training
LLaVA-Cot-11B (Xu et al., 2025) GPT-4o 100k
Mulberry-7B (Yao et al., 2024) GPT-4o, Qwen2-VL-72B 260k
Vision-R1-7B (Huang et al., 2025) Deepseek-R1 200k + 10k
OpenVLThinker-7B (Deng et al., 2025) DeepSeek-R1-Distill-Qwen-14B 59.2k
MM-EUREKA-Qwen-7B (Meng et al., 2025) 15k
@@
After conducting MCTS-based sample selection and obtaining a filtered set of high-quality training data (11k for 7B and 7.5k for 72B), we then perform RL fine-tuning on the Qwen2.5-VL models using these selected data. Specifically, we employ Group Relative Policy Optimization (GRPO) loss function proposed by (Shao et al., 2024b) for training, with the objective defined as follows:
@@ 4.1 Benchmark Evaluation
• For 7b-level VLMs, we use Qwen2.5-VL-7B-Instruct as the base model and perform RFT on the 11k high-quality data obtained through MCTS-based filtration, resulting in our reasoning model, named ThinkLite-VL-7B. We conduct training using Easy-R1 (Zheng et al., 2025) code base and set GRPO rollout number as 32. Our main baselines are as follows: (1) Qwen2.5-VL-7B-Instruct (Bai et al., 2025), serving as our base model; (2) ThinkLite-VL-Random11k, trained using RFT on a randomly sampled subset of 11k instances from the full 70k dataset. Besides, we report the performance of several recent general and reasoning VLMs for comparison, including general opensourced models LLaVA-Onevision-7B (Li et al., 2024a) and InternVL2.5-8B (Chen et al., 2024e), the SFT-based reasoning models LLaVA-Cot-11B (Xu et al., 2025) and Mulberry-7B (Yao et al., 2024), as well as the RFT-based reasoning models Vision-R1 (Huang et al., 2025), MM-Eureka-Qwen-7B (Meng et al., 2025), and OpenVLThinker-7B (Deng et al., 2025).
• For 72B-level VLMs, we use Qwen2.5-VL-72B-Instruct as the base model. We perform RFT on the 7.5k high-quality data obtained by Qwen2.5-VL-72B-Instruct through MCTS-based filtration and get 72B reasoning model ThinkLite-VL-72B. The 72B-level baselines include: (1) our base model Qwen2.5-VL-72B-Instruct (Bai et al., 2025); (2) two opensourced general VLMs LLaVA-Onevision-72B (Li et al., 2024a) and InternVL2.5-78B (Chen et al., 2024e); (3) one opensouced reasoning model QvQ-72B (Wang et al., 2024b); (4) ThinkLite-VL-Random7.5k, trained using RFT on 7.5k randomly selected samples from the full 70k dataset. We also include proprietary models as performance references which include OpenAI-GPT-4o and OpenAI-o1. For all models, we use 8$\times$80G A100 GPUs for model training and evaluation.
@@
We select eight widely used VLM benchmarks for evaluation, namely MathVista (Lu et al., 2024), MathVison (Wang et al., 2024a), MathVerse (Zhang et al., 2024a), MMMU (Yue et al., 2024), MMStar (Chen et al., 2024c), MMBench (Liu et al., 2024c), MMVet (Yu et al., 2024b), and AI2D (Kembhavi et al., 2016). Among them, MathVista, MathVison, and MathVerse are widely used in VLM research to evaluate mathematical reasoning capabilities, while MMVet also includes a significant number of mathematical reasoning tasks. In contrast, MMMU, MMStar, MMBench, and AI2D are primarily utilized to assess VLM’s visual perception reasoning and scientific reasoning abilities.
Table 2: Comparison of different VLMs on 8 widely used visual benchmarks. Our model achieves SoTA performance at both 7B level and 72B level on 6 benchmarks and reaches a SoTA performance of 79.7 on MathVista among all VLMs. On average, our model improves performance by 7.5% and 6.5% compared with our base models Qwen2.5-VL-7B-Instruct and Qwen2.5-VL-72B-Instruct. We do not evaluate Mulberry-7B on MathVision because Mulberry-7B uses MathVision as training dataset. We evaluate all models with same code using vLLM (Kwon et al., 2023) inference. For reasoning models, we use thinking templates provided in their codebase to generate thoughts and get the final answer.